3. The electric (E) and magnetic (H) field vectors in a medium are given by


𝑬 = π’Š^ 20π‘π‘œπ‘ (πœ”π‘‘ + πœ‹π‘¦)𝑉/π‘š , 𝑯 = π’Œ^ 20 π‘π‘œπ‘ (πœ”π‘‘ + πœ‹π‘¦)𝐴/π‘š


πœ‚


Use Maxwell’s equations to determine Ξ· and Ο‰. Assuming that for the medium πœ€ = 36πœ€π‘œ ,


πœ‡ = πœ‡π‘œ .

1
Expert's answer
2022-03-24T18:13:15-0400

"\eta=\frac EH=1,"

"\omega=\frac c{\sqrt{\varepsilon \varepsilon_0\mu\mu_0}}=5\cdot 10^7."